OpenStudy (anonymous):

Explain how to do this Lim 3x^2(x+1) /(over) (x +1) x->-1 = ?

OpenStudy (saifoo.khan):

Hey matt. :D

myininaya (myininaya):

\[\lim_{x \rightarrow -1}3x^2=3(-1)^2=3(1)=3\]

OpenStudy (anonymous):

\[ \frac{x+1}{x+1}=1\]

OpenStudy (anonymous):

Hey !

OpenStudy (anonymous):

How would you calculate that limit using the Definition?

OpenStudy (anonymous):

let me rephrase that \[\frac{x+1}{x+1}=0 , x\neq -1\]and since we are taking a limit as x approaches -1 we know for certain \[x\neq -1\]

myininaya (myininaya):

=0?

OpenStudy (anonymous):

TYPO!!

myininaya (myininaya):

let me rephrase what sallielte rephrased

OpenStudy (anonymous):

See Attachment ^ .

myininaya (myininaya):

\[\frac{x+1}{x+1}=1 , x\neq -1 \]

myininaya (myininaya):

any questions matt?

OpenStudy (anonymous):

Kinda makes sense